### v1

# Experimental Study of Bremsstrahlung Gamma Ray Emission and Short-Range Correlations in 124Sn+124Sn Collisions at 25 MeV/u

Abstract
Short-range correlation (SRC) in nuclei refers to the nucleons forming temporally correlated pairs in close proximity, giving rise to the high momentum of the nucleons beyond the Fermi surface. It has been reported that bremsstrahlung $\gamma$ production from neutron-proton process in heavy-ion reactions provides a potential probe of the existence of SRC in nuclei. In this paper, we present in detail the precision measurement of bremsstrahlung $\gamma$-rays in $\rm{}^{124}Sn$$\rm{}^{124}Sn$ reactions at 25 MeV/u using the Compact Spectrometer for Heavy IoN Experiment (CSHINE). The experimental setup, detector calibration, trigger scheme and data analysis procedures as well as the model comparison are presented in detail. Background contributions are subtracted using two methods to ensure robustness. By comparing the experimental $\gamma$ spectrum with the Isospin-dependent Boltzmann-Uehling-Uhlenbeck simulations, the high momentum tail (HMT) fraction of $R_{\rm HMT}=(20\pm 3)\%$ is derived in 124Sn nuclei. This work presents the detailed experimental measurement and analysis framework for the precise determination of the HMT fraction via bremsstrahlung $\gamma$-ray emission, demonstrating a new paradigm to study nucleon SRCs in nuclei using low-energy heavy-ion collisions.
I Introduction
Short-range correlation (SRC) [1, 2, 3, 4] of nucleons is a temporal fluctuation that nucleons form temporally correlated pairs in close proximity in atomic nuclei. SRC plays an essential role in understanding the nuclear structure as well as to shed insight to the properties of dense nucleonic matter. Most traditional nuclear models are based on mean-field theory, in which nucleons are confined below the Fermi momentum and quark-level interactions are largely neglected. However, the EMC effect [5] indicates that deviations in nucleon momentum distributions from shell model predictions stem from the underlying quark structure of nucleons [6, 7, 8, 9, 10]. This is exemplified by the fact that nuclear parton distribution functions (PDFs) in heavy nuclei deviate markedly from the naive sum of those in free nucleons [11, 12, 13, 14, 15, 16], underscoring the importance of understanding the momentum distributions of quarks and gluons within bound nucleons and nuclei. SRCs, a ubiquitous feature of all nuclei, originate from fluctuations in the nuclear ground state and predominantly involve neutron-proton ($np$) pairs [17, 3, 18, 19, 20, 21, 22], driven by the strong tensor component of the nucleon-nucleon interaction at sub-Fermi-range distances [23, 24, 25]. Notably, $np$ pairs often display significant spatial overlap, which may favor isosinglet, spin-singlet $ud$ diquark configurations at the quark level [22]. These pairs also exhibit large relative momenta, contributing to the emergence of a high momentum tail (HMT) in the nucleon momentum distribution beyond the Fermi surface [19, 26]. Since SRCs reflect dynamics beyond the reach of mean-field approximations, incorporating both short-range nuclear forces and potential quark-level effects [27], they have become a focal point of contemporary theoretical and experimental research in nuclear physics.
Electron scattering experiments have proven to be particularly powerful in probing SRCs, owing to the well-understood nature of electromagnetic interactions. By carefully selecting kinematic conditions, complex final-state effects can be minimized [28], thereby improving the sensitivity to SRCs. In exclusive measurements [29, 3, 18, 30, 19], both the scattered electron and the two nucleons from the correlated pair are detected. A high-energy electron with large momentum transfer is used to knock out one nucleon from the SRC pair. Strong back-to-back correlations [1] are observed in knocked-out nucleon pairs with momenta above the Fermi momentum, whereas no angular correlation is seen below it. In inclusive quasielastic (QE) electron scattering [31, 32, 33, 34, 35, 36], only the scattering electrons are measured and the momentum distribution of nucleons inside the nucleus can be probed. In particular, the scattering experiments related to low-energy transfer side of the QE peak can provide evidence for nucleons with high momenta [35, 37], indicating the presence of short-range correlated pairs.
Hadronic reactions also serve as one of the primary approaches in the SRCs experimental studies. In exclusive scattering experiments involving large momentum transfer proton-nucleus collisions [38, 17], the momentum of the struck proton can be reconstructed while simultaneously measuring the momentum of the correlated neutron. Under high momentum transfer conditions, the struck fast-bound proton can be described using the instantaneous approximation, which significantly enhances the resolution of nuclear structure. Moreover, high-energy inverse kinematics scattering [39], where an incoming ion beam collides with a proton target, has emerged as a novel technique to investigate SRCs. By detecting the knocked-out protons and residual nuclear fragments, this method effectively suppresses complications from initial- and final-state interactions between hadrons [40, 41], thereby improving the reliability of reconstructed particle distributions.
Heavy-ion collisions (HICs) provide the only means of producing and studying extreme states of nuclear matter in terrestrial laboratories. In the 1980s, high-energy photon emission observed in HICs emerged as a puzzling phenomenon [42, 43, 44], drawing growing interest due to its unclear origin. Subsequent experimental measurements have shown that a significant fraction of these $\gamma$-rays originate from bremsstrahlung processes induced by $np$ collisions in the early-stage of HICs [45, 46, 47, 48, 49, 50, 51, 52]. Recent advances in the study of SRCs have substantially deepened our understanding of nuclear structure. Theoretical studies indicate that high-energy bremsstrahlung $\gamma$-ray production in HICs [53] is sensitive to the high-momentum components of nucleon momentum distributions generated by SRCs [54], rendering hard photons in the $\gamma$-ray spectrum a promising probe of these correlations [53, 55]. Notably, photons interact only weakly with the nuclear medium and are largely unaffected by final-state interactions, making them a particularly clean and sensitive probe of short-range nuclear dynamics. Using the isospin-dependent Boltzmann-Uehling-Uhlenbeck (IBUU) transport model [56] with momentum-dependent interactions (MDI) [57], one can simulate $\gamma$-ray spectra from HICs by varying the fraction of high-momentum nucleons in the initial nuclear state. Comparing these simulations with experimental bremsstrahlung data enables the extraction of information on SRCs in nuclei.
To explore the underlying dynamical mechanisms in intermediate-energy HICs, the Compact Spectrometer for Heavy Ion Experiment (CSHINE) [58, 59, 60] has been constructed, maintained, and continuously upgraded at the Radioactive Ion Beam Line in Lanzhou (RIBLL-1). Through multiple beam campaigns, CSHINE has developed advanced detection technologies and delivered a wide range of physics results in key areas including the equation of states of nuclear matter, isospin evolutions [61, 62, 63], short-range correlations in nuclei [64, 65, 66], neutron-neutron interactions [67], and the interplay between fission dynamics and isospin dynamics [68]. To date, CSHINE has completed five beam campaigns. As part of the detection system, a CsI(Tl)-based total absorption $\gamma$-ray spectrometer array (CSHINE-Gamma) [69] was first deployed during the fourth experiment in 2022, which involved 25 MeV/u 86Kr + 124Sn reactions. In this experiment, CSHINE-Gamma successfully detected and analyzed $\gamma$-ray spectra emitted from HICs. The data revealed an nonzero HMT fraction, providing experimental indication for the presence of SRC-induced HMT components in nuclei at about 90% confidence level [66, 65]. However, this early study faced limitations due to the restricted energy coverage of CSHINE-Gamma and a narrow event reconstruction region confined to the central scintillator. These constraints hindered the ability to quantify the HMT fraction with high statistical confidence. Moreover, due to the use of an asymmetric reaction system, it was not possible to determine the SRC fraction for a specific nucleus.
In 2024, the fifth CSHINE experiment, involving a 25 MeV/u $\rm{}^{124}Sn$$\rm{}^{124}Sn$  beam, introduced significant upgrades of CSHINE-Gamma specifically aimed at enhancing high-energy $\gamma$-ray detection. These improvements included an extended energy range, expanded event reconstruction regions, and overall enhanced detection efficiency of high-energy $\gamma$-rays. As a result, the statistics for high-energy bremsstrahlung $\gamma$-rays were significantly increased. Leveraging the unprecedented statistical precision and improved measurement accuracy of this experiment, the HMT fraction in the nucleon momentum distribution of 124Sn nuclei has been determined quantitatively with improved precision [64]. This result marks the first high-precision extraction of the SRC fraction in low-energy HICs, achieved using bremsstrahlung $\gamma$-rays as a clean and unambiguous probe.
In this paper, we present the analysis and results of the 25 MeV/u $\rm{}^{124}Sn$$\rm{}^{124}Sn$  experiment, providing in detail the technical underpingnings of a submitted letter [64]. Specifically, we describe the detector setup, calibration and data analysis procedure. The final bremsstrahlung $\gamma$ spectra were derived using different methods, all of which were compared with the transport model simulations, yielding a consistent conclusion on the fraction of SRC in 124Sn nuclei. This paper is organized as following: Section II outlines the experimental setup, including the CSHINE detectors and $\gamma$ hodoscope calibration. Section III details the data analysis procedure. Section IV describes the IBUU-MDI model and parameter tests theoretically. Section V presents results and discussion, focusing on data uncertainties and HMT fraction analysis of measured $\gamma$ spectra. Section VI reconstructs the original $\gamma$ spectrum using Richardson-Lucy deblurring method and compares it with IBUU-MDI predictions directly. Section VII concludes with a summary and outlook.
II Experimental Setup
II.1 Detector Configuration of CSHINE
The experiment was performed at the Radioactive Ion Beam Line at Lanzhou (RIBLL-1). The beam of $\rm{}^{124}Sn$  at 25 MeV/u was delivered by the Heavy Ion Research Facility at Lanzhou (HIRFL) and bombarded on a $\rm{}^{124}Sn$  target of 1 $\rm mg/cm^{2}$ installed in the chamber located at the final focal plane of RIBLL-1. The products of the $\rm{}^{124}Sn$$\rm{}^{124}Sn$  reactions were measured by the Compact Spectrometer for Heavy Ion Experiment (CSHINE), which is dedicated to the studies of heavy ion reactions at Fermi energies. The schematic view of the CSHINE setup in this experiment is presented in Fig. 1. The light charged particles (LCPs) were detected by 8 silicon-strip detector telescopes (SSDTs), six of which provided good particle identification (PID). Each SSDT consists of a single-sided silicon strip detector (SSSD), a double-sided silicon strip detector (DSSD) and a $3\times 3$ CsI(Tl) array. The overall energy resolution of each SSDT is better than 2% [59], the pixel size of the SSDT is $4\times 4$ mm. The charged isotopes for $Z\leq 6$ can be clearly identified. The performance of the SSDTs had been demonstrated in previous experiments, one can refer to [58, 59, 70] for details.
Figure 1: (Color Online) Experimental setup of CSHINE. The $\gamma$ hodoscope is located at $\theta_{\rm lab}=110^{\circ}$ to measure the bremsstrahlung $\gamma$-rays from the collisions of the 124Sn projectile on the 124Sn target.
Figure 2: (Color Online) Trigger scheme of the $\rm{}^{124}Sn$$\rm{}^{124}Sn$  beam experiment.
The neutrons are measured by a plastic scintillator array consisting of $4\times 5$ units, covering the laboratory polar angle $17^{\circ}<\theta_{\rm lab}<53^{\circ}$ in partial azimuth. Each unit is a block of plastic scintillator with the size of $\rm 15\times 15\times 15~cm^{3}$ readout by photomultiplier tube (PMT). The distance of the neutron array to the target is 200 cm. The neutron energies are measured by the time of flight (TOF) method. The start timing of TOF is provided by 4 $\rm BaF_{2}$ fast scintillators surrounding the target fired by the reaction $\gamma$-rays. For the performance of the neutron array, one can refer to [71]. Besides the neutron wall, a liquid scintillator neutron detector, with the sensitive volume of $\rm\phi 34.8~cm\times 102~cm$, was installed at 513 cm to the target at $\theta_{\rm lab}=60^{\circ}$. Two 8-inch PMTs are attached to the left and right ends of the sensitive volume.
The bremsstrahlung $\gamma$-rays of this analysis are measured by a CsI(Tl) hodoscope (CSHINE-Gamma), containing 15 units in a $4\times 4$ configuration (with one corner being vacant). Each CsI(Tl) unit has the dimension of $\rm 7\times 7\times 25~cm^{3}$ and is coupled to a PMT of Hamamatsu R2631 for signal readout. The radiation length of CsI(Tl) is $X_{0}=8.39\rm~g/cm^{2}$ and the Moliere radius is $R_{\rm M}=3.531~\rm cm$. The energy resolution of the units is about $3.6\%$ for 1 MeV $\gamma$-ray and about 2% for $\gamma$-ray energy beyond 10 MeV. In order to suppress the cosmic-ray muons, 3 thick plastic scintillators ($5\times 30\times 30\rm~cm^{3}$) were mounted surrounding the CsI(Tl) hodoscope on top, left and right sides. The linearity of the CsI(Tl) response to high-energy $\gamma$-rays was tested after the experiment using the quasi monochromatic $\gamma$ beam at the Shanghai Laser Electron Gamma Source (SLEGS) [72]. The performance of the $\gamma$ hodoscope can be found in [69].
In order to determine the centrality event-by-event, a small CsI(Tl) array was installed at the forward angle in the reaction chamber. The array are split into two modules and placed on the left and right side of the beam, containing $4\times 8$ units, respectively. Each unit has the dimension of $1\times 5\rm~cm^{3}$ and coupled to a PMT. This small CsI(Tl) array provides the multiplicity information of the charged particles (The $\gamma$-rays leaves much lower energy deposit compared to the charged particles). This small CsI(Tl) is not included in the trigger scheme.
II.2 Trigger Scheme
Fig. 2 presents the trigger scheme in the $\rm{}^{124}Sn$$\rm{}^{124}Sn$  experiment. The trigger scheme of the experiment is constructed using Field Programmable Gate Array technology [73], which has been integrated in the CAEN V2495 module. The timing signals (of NIM standard) from different detectors are processed by the firmware written in advance in the V2495 module for further logical operations. The timing information of the CsI(Tl) $\gamma$ hodoscope is provided by leading edge discriminator integrated in the N914 module. The timing signal of the neutron wall, $\rm BaF_{2}$ and liquid scintillator neutron detector are extracted from the front end electronics (FEE) and sent to constant fraction module (CF8000) to generate a NIM signal of timing. These NIM signals are further split into two branches, one goes to TDC and the other to the trigger diagram. The time of the SSDTs, however, were extracted from the TOUT of the main amplifier (MSCF16). A TRIG signal on the front panel of MSCF16 is added (via OR operation) to form the SSD M1  trigger condition (see below), while the Multilicity signal from the rear panel of each MSCF16 is extracted and summed up to generate an analog signal with the amplitude being $50M$ mV, where $M$ is the multiplicity. This signal is discriminated by a CF8000 to generate a SSD M2  trigger condition (see below). The time walk of this trigger condition is very large and requires careful treat. In total, six trigger conditions are generated as listed below. The final trigger signals will be generated if any of the 6 trigger conditions is satisfied.
Trig1: SSD M2, Two (or more) LCPs are detected in SSDTs.
Trig2: SSD M1 & CsI M1, One (or more) LCP is detected in SSDTs (SSD M1) and one (or more) unit of the high energy $\gamma$ hodoscope fires (CsI M1).
Trig3: SSD M1 & NA M1, One (or more) LCP is detected in SSDTs (SSD M1) and one (or more) unit of the neutron array fires (NA M1).
Trig4: SSD M1 & LS, One (or more) LCP is detected in SSDTs (SSD M1) and the liquid scintillator is fired at both ends.
Trig5: NA M1 & T0, One (or more) unit of the neutron array fires (NA M1) and one (or more) $\rm BaF_{2}$ ($T_{0}$) is fired.
Trig6: LS & T0, The liquid scintillator is fired in both readout PMT and one (or more) $\rm BaF_{2}$ ($T_{0}$) is fired.
Through the monitoring endpoints, the NIM signals of all the 6 predefined trigger conditions are sent to a TDC channel for saving. Thus, one can decode the trigger type event-by-event from the self-triggered peak on the corresponding TDC spectrum.
II.3 Calibration of the $\gamma$ Hodoscope
Since our goal is to measure the high energy $\gamma$-rays of tens of MeV, while the radioactive $\gamma$ source usually has the energy of a few MeV, one must be careful to conduct the calibration and evaluate the uncertainty originated from the calibration.
The amplitude signal of each CsI(Tl) unit was extracted from both ADC-E ($\times 1$ gain) and ADC-XE ($\times 10$ gain) to record the high energy and the low energy $\gamma$-rays, respectively. This enables to utilize the dual-range method to do the energy calibration in three steps.
Step 1: Calibrate the ADC-XE channel using radioactive $\gamma$ source of 60Co and 232Th, emitting $\gamma$-rays at 1.17, 1.33 and 2.61 MeV, respectively. Sufficiently good linearity in the $\gamma$ energy range of the source can be assured [69]. Using a linear fitting, the $\gamma$ energy in MeV varying with the ADC-XE output in channel number $\rm{CH_{XE}}$ can be written as:
$E_{\gamma}=a\text{CH}_{\text{XE}}+b$ (1)
Step 2: Build the linear relationship between ADC-E and ADC-XE by fitting the correlation of the ADC-E and the ADC-XE output in a wide range. The relation can be written as
$\text{CH}_{\text{E}}=\alpha\text{CH}_{\text{XE}}+\beta$ (2)
Step 3: Perform the linear calibration of ADC-E. According to the Eq. (1) and Eq. (2), one can transform the linear calibration of ADC-XE to the ADC-E channel. Substituting Eq. (1) to Eq. (2), one writes
$E_{\gamma}=\frac{a}{\alpha}\text{CH}_{\text{E}}+\frac{b\alpha-a\beta}{\alpha}.$ (3)
Fig.3 illustrates the calibration procedure using unit CsI05 as an example. Panel (a) presents the linear calibration of ADC-XE using $\gamma$-rays from radioactive sources. The three data points correspond to the 1.17 MeV and 1.33 MeV $\gamma$-rays from 60Co and the 2.61 MeV $\gamma$-rays from 232Th, respectively. The red line denotes the linear fitting. Panel (b) displays a two-dimension histogram of ADC-E versus ADC-XE obtained from online data, with the red line representing the linear fit. A clear linear correlation is observed over a wide range. With these two linear fits, the relationship between energy and ADC-E can be established by combining the ADC-XE and ADC-E correlation with the ADC-XE and energy calibration.
The dual-range method calibration has been verified after the beam experiment at SLEGS, using the quasi monochromatic $\gamma$ beam up to 17 MeV [72]. Clearly, the final question remaining on the calibration is the non-linearity of the CsI(Tl) response to high energy $\gamma$-rays. According to the test results at SLEGS, the non-linearity is at the level to 2% to 4% [72]. The variation of the energy spectra caused by the linear calibration is taken as one origin of the systematic uncertainty.
Figure 3: (Color Online) The dual-range calibration procedure (Taking CsI05 for example). (a) The calibration of ADC-XE using radioactive source. (b) The linear relationship between the ADC-E and ADC-XE channels.
During the beam experiment, the source calibration has been done for 3 times in order to monitor the possible shift of the PMT output with time due to the change of environment. It is found that the shift of the PMT amplification is insignificant and the variation is treated as one origin of the systematic uncertainty as well. Table 1 presents the set of the calibration coefficients corresponding to the highest statistics of data, which are used to calibrate the central energy spectrum.
Table 1: The calibration coefficients of the 15 CsI(Tl) units.
CsI Unit $\alpha$ $\beta$ $a$ $b$
0 0.099764(5) 102.53(1) 0.00290(27) -0.16(14)
1 0.099738(9) 64.162(14) 0.00301(17) -0.07(10)
2 0.099852(9) 92.780(14) 0.00296(17) -0.15(9)
3 0.099795(8) 49.083(14) 0.00279(13) -0.08(8)
4 0.100050(8) 81.447(13) 0.00305(15) -0.18(8)
5 0.099657(10) 60.506(16) 0.00305(16) -0.03(8)
6 0.09982(1) 75.060(15) 0.00312(15) -0.11(8)
7 0.099741(7) 65.341(12) 0.00296(16) -0.02(8)
8 0.100038(8) 91.602(14) 0.00298(15) -0.14(8)
9 0.100137(9) 60.988(15) 0.00299(16) -0.05(8)
10 0.099849(9) 69.493(15) 0.00287(13) -0.04(7)
11 0.099877(9) 77.104(14) 0.00297(18) -0.05(10)
12 0.099383(9) 83.566(15) 0.00286(22) -0.17(12)
13 0.099900(8) 63.517(13) 0.00304(15) -0.07(8)
14 0.099691(9) 106.817(15) 0.00298(16) -0.13(9)
III Data analysis Procedure
In this section, we focus on the data analysis of the CsI(Tl) hodoscope, along with the performance of the detector.
III.1 Time-Amplitude Correction of Time Signals
[…]
Due to the nature of leading-edge discrimination by the N914 module, and given that the rising time of the CsI(Tl) is at the order of $10^{2}$ ns, the timing signal exhibits a signal amplitude-dependent shift (time walk). Thus, a time–amplitude (T–A) correction is necessary.
Fig. 4(a) shows the correlation between the ADC-E (energy) and timing signals for a unit (unit CsI05 for example) of the high energy $\gamma$ hodoscope. Here the timing signal in unit of ns is defined as the time difference between the crystal unit and the BaF2 start-time detector. Thus, the data points on the scattering plot merely correspond to the events with the $\rm BaF_{2}$ being fired coincidentally. Inspecting the scattering plot carefully, one finds two interesting features. 1) For the signals with large amplitude, the timing distribution is quite sharp, indicating the trigger signal is generated by the BaF2 detector. For the low energy part, on the other hand, the time distribution is very broad, indicating the contributions of the random coincidence between CsI(Tl) and $\rm BaF_{2}$ caused by the residual $\gamma$-rays. 2) For the events near the trigger peak, a clear time-amplitude (T-A) correlation appears, as guided by the red curve. This is caused by time walk and shall be corrected. Although this effect does not significantly impact the online acquisition or energy calibration, it does affect the time correlation between different detector units, thereby compromising the reconstruction accuracy of single-photon events.
Figure 4: (Color Online) Time distribution and correction. (a) Correlation between ADC-E energy signals and timing signals ($t_{\text{det}}$) for a gamma detector unit (unit CsI05 for example), along with reciprocal function fitting (red lines). (b) Correlation between ADC-E energy signals and corrected timing signals for a gamma detector unit (unit CsI05 for example). Here, both $t_{\text{det}}^{\prime}$ and $t_{\gamma}^{\prime}$ have been offset by the timing of the BaF2 detector.
To improve time correlation precision and eliminate the systematic error introduced by amplitude-dependent time shifts, we apply a reciprocal functional form to fit the correlation between ADC-E and timing signals. The incident $\gamma$-ray time $t_{\gamma}$, is expressed as:
$t_{\gamma}=t_{\text{det}}-t_{\text{corr}}(\text{CH}_{\text{E}}),$ (4)
where $t_{\text{det}}$ is the TDC channel converted to nanoseconds, and $t_{\text{corr}}(\text{CH}_{\text{E}})$ represents the T–A correction term, modeled as a function of the signal amplitude from the same crystal unit. Assuming the correction is inversely proportional to amplitude, it can be expressed as:
$t_{\text{corr}}(\text{CH}_{\text{E}})=\frac{C_{0}}{E_{0}-\text{CH}_{\text{E}}}+t_{\text{const}}.$ (5)
Here, $C_{0}$ is a time constant, and $E_{0}$ is the baseline offset for that specific detector unit. After applying the T–A correction, the resulting corrected time $t_{\text{const}}$, which reflects the actual $\gamma$ emission time, becomes independent of signal amplitude. Since the photon flight time is fixed, $t_{\text{const}}$ can also be treated as a measurable constant for a given type of physics event.
To calibrate the three constants involved, we manually selected more than five hotspots for each unit and performed fitting using the reciprocal function. The red lines in the plot of Fig. 4(a) illustrate the fitted results for the representative unit. As shown in Fig. 4(b), taking unit CsI05 as an example, after applying the time correction, the T-A correlation of the $\gamma$-ray events in the region near the trigger peak has been effectively corrected, resulting in a vertical band, indicating that the time is no longer dependent on the signal amplitude. Meanwhile, the trigger peak has been aligned with the reference time from the BaF2 start-time detector and is used as the time zero point. The fitted parameters for all CsI(Tl) units of the $\gamma$ hodoscope are summarized in Table. 2.
Table 2: Time correction parameters for each unit in the bremsstrahlung $\gamma$ detector array.
CsI Unit $C_{0}$ $E_{0}$ $t_{\text{const}}$
0 19190 -381.198 592.276
1 12542.3 -218.541 588.291
2 16874.5 -443.026 577.473
3 18759.2 -607.21 592.099
4 17320 -495.872 592.945
5 15157.8 -258.436 591.224
6 10046.1 -193.933 575.274
7 13821.7 -400.981 584.002
8 11020.5 -291.473 573.071
9 13329.8 -310.908 584.707
10 10879 -235.12 586.769
11 9897.46 -295.675 582.07
12 7312.49 -67.4275 583.091
13 9322.57 -127.872 586.164
14 14622.8 -272.735 594.726
These calibration parameters were subsequently applied to correct the timing signals in the experimental data, and all time information used in the following analyses has been corrected using these parameters derived by the above method.
III.2 Temporal coincidence of neighbouring crystals
[…]
Fig. 5 (a) presents the timing correlation between two neighboring CsI(Tl) units, taking CsI05 and CsI06 for an example. A band following the trend of $\rm T_{5}\approx T_{6}$ exists, marking the timing coincidence between these two units for these events. In addition, considerable portion of events are found far away from the band, indicating the coincidence is lost. A sharp peak can be found in the vicinity of $270$ ns along $\rm T_{5}$ and $\rm T_{6}$. It corresponds to the self-triggered peak, where the scintillator unit itself initiates the trigger. Near $T\approx 150$ ns, the distribution exhibits a sudden decrease in rate. The time difference between 150 and 270 ns reflects approximately the width of the trigger coincidence window. For the events far away from the $\rm T_{5}\approx T_{6}$ band, the two detectors do not coincide in time, and shall be excluded for high energy $\gamma$ reconstructed. As expected, if we select the high energy events with a cut condition $E_{5}+E_{6}\geq 30$ MeV, the correlated events situate dominantly in the vicinity of $T_{5}\approx T_{6}$, as shown in Fig. 5 (b).
Figure 5: (Color Online) Time correlation between two neighbouring CsI(Tl) crystal, No. 5 and No. 6, Without energy cut (a) and with energy cut of $E_{5}+E_{6}\geq 30$ MeV (b).
Fig. 6 (a) and (b) shows further the time distribution of unit CsI05 and the time difference between the two crystals referred in Fig. 5, respectively. Again, the sharp peak corresponds to self-triggered gamma events. The main peak extending about 120 ns on the right side of the self-triggered peak represents passively triggered $\gamma$ events, which are recorded in response to the trigger signal determined temporally by other detectors. The extended tail in the time distribution arises from other events, also passively recorded, of which the timing signal stays far from the main trigger windows because of two possible reasons, the recorded $\gamma$-ray originates from a randomly coincidence event, or the $\gamma$-ray is a truly coincidence event, but the trigger timing is not well determined (see the subsection of III.5). The time difference between unit CsI05 and CsI06, as shown in Fig. 6 (b), exhibits a narrow peak. To ensure that the detected signals originate from the same physical event, a timing cut of $|\Delta T|\leq 50~\text{ns}$ between two neighboring units is applied during the $\gamma$-ray energy reconstruction process.
Figure 6: (a) Time distribution of unit 5. (b) Time difference between two neighboring crystals (CsI05 and CsI06 for example).
III.3 Energy Reconstruction Algorithm
Based on the configuration of the detector array and the characteristic features of $\gamma$-ray events emitted from HICs, we developed a dedicated event reconstruction algorithm that reliably aggregates energy deposits from different CsI crystal units belonging to the same physical event.
Fig. 7 illustrates the spatial scheme considered in the event reconstruction algorithm. The $4\times 4$ grid represents the layout of the CsI scintillator array, with each cell labeled according to its detector unit number in the view facing the front surface of the hodoscope. Panel (a) presents a color-coded classification of the array, highlighting the 4 central units, 8 edge units, and 3 corner units. The plastic veto detector array is positioned on three sides of the CsI array, i.e., left, right and top. Due to the absence of a veto detector on the bottom side, units 1 and 2 are excluded from edge-type reconstruction. For events in which a event core with the largest energy deposit is identified (see next paragraph), the reconstruction algorithm sums the energy depositions from all neighboring detectors within a $3\times 3$ area centered on the core unit. As illustrated in panel (b), when unit 10 is the event core, energy contributions from units 5, 6, 7, 9, 11, 12, 13, and 14 are included to recover potential leakage. For edge units, shown in panel (c), the energy of five adjacent detectors is incorporated into the reconstruction only if no signal is detected in the corresponding veto scintillators. Taking unit 7 as an example, the algorithm adds energy from units 2, 3, 6, 10, and 11.
Figure 7: Schematic of the event reconstruction algorithm. (a) Layout of the CsI array with detector unit numbers. (b) Reconstruction for central units (example: unit 10). (c) Reconstruction for edge units (example: unit 7).
The basic principle of the reconstruction algorithm involves firstly determining the “core” of the event and then using temporal coincidence to identify its “periphery” before summing up the leakage energy. The reconstruction algorithm proceeds as follows:
First, identifying the event core. The 15 CsI detector units are sorted in descending order according to their energy deposit. If the highest energy deposit is below the energy threshold (1 MeV), the algorithm immediately returns an empty result. The sorted units are then traversed sequentially to identify “cores” for reconstruction, subject to the following criteria:
  • The deposited energy of the candidate detector unit must exceed the threshold (1 MeV);
  • The time difference between the candidate detector unit and any existing center must be sufficiently large (greater than 100 ns) to ensure separation of distinct physical events;
  • If the time difference is small (less than 100 ns), a further check is performed: if the candidate lies within the reconstruction range of an existing center and the time difference is smaller than the specified limit (50 ns), the candidate is rejected and can not be a new event core.
Second, Calculating the total energy. Once a core is identified, the deposited energies of neighboring detectors within the defined spatial and temporal criteria are aggregated to complete the reconstruction of the current event.
If the “core” lies in the central 4 blocks of the $\gamma$ array detector, the reconstruction energy of the “core” is calculated as the sum of the energy deposited in the core and that in its $3\times 3$ neighboring blocks, provided the time difference between the core and the neighboring blocks is within 50 ns. If the “core” is located on the edge of the $\gamma$ hodoscope and there is no signal in the adjacent plastic scintillator veto detector, it indicates that the $\gamma$ photon has not escaped and shall be recorded. In this case, the energy deposited in neighboring scintillators meeting the temporal coincidence criteria is also included in the core’s reconstructed energy.
With each triggered event the data in all units in the CSHINE-Gamma array are recorded. It is important to check the multiplicity of the fired units in each $\gamma$ incident event and the number of temporally separated $\gamma$ events for each trigger. Fig. 8 (a) presents the correlation between the multiplicity of fired CsI(Tl) units and the total reconstructed energy for the single $\gamma$ event. It can be seen that events with lower total energy generally involve fewer reconstructed CsI units. For the events with very high energy ($E_{\rm tot}>100$ MeV), they are mainly the cosmic ray muons, for which the average multiplicity is high. For the events with low energy ($E_{\rm tot}<20$ MeV), we must bring here the caution that the multiple fired CsI(Tl) units do not necessarily record one single incident $\gamma$-ray. Instead, they may correspond to multiple $\gamma$-rays with even lower energy. But anyway the low energy part of the spectrum is not of the interest of analyzing the bremsstrahlung $\gamma$-ray emission.
Figure 8: (Color Online) (a) Cluster size (number of reconstructed CsI units) as a function of reconstructed energy $E_{\gamma}$) (b) Number of event core $N_{\rm core}$ per triggered event with (red histogram) and without (blue histogram) the cut of $E_{\rm tot}>35$ MeV.
Fig. 8 (b) displays the distribution of the number of reconstructed cores per trigger (blue histogram). It is observed that except for $N_{\text{core}}=1$ cases, there are about 10% contributions from $N_{\text{core}}>1$, indicating that multiple reactions are recorded following one triggered event. Applying an additional condition of $E_{\rm tot}>35$ MeV (red histogram), however, the result reveals that the $N_{\rm core}$ distribution is overwhelmingly dominated by $N_{\rm core}=1$, and the contribution of $N_{\rm core}=2$ are reduced to about 0.1%. It is clearly demonstrated that only one high energy bremsstrahlung $\gamma$-ray is reconstructed for each triggered event.
III.4 Feature of the reconstructed $\gamma$ Events
Based on the event reconstruction algorithm described above, we are able to count the number of detected $\gamma$-rays as a function of energy, thereby generating the corresponding $\gamma$-ray energy spectrum.
To examine the energy spectra of individual CsI detector blocks as well as the spectra reconstructed using the event reconstruction algorithm (considering the 4 central and 6 edge units), we performed a block-by-block analysis. Fig. 9 presents the experimental $\gamma$-ray energy spectra obtained in coincidence with reaction on the target, using the reconstruction method and selection criteria described above. The blue histograms represent the original energy spectra recorded by each individual CsI block, while the red histograms show the energy deposit in the corresponding unit after event reconstruction for the 4 central and 6 edge blocks. On the spectra of individual blocks, there is a bump at about 60 MeV, which corresponds the cosmic ray mouns. After the reconstruction procedure, the spectrum on each single unit exhibits a relatively clean and smooth descending trend, and shows similar shape in all units.
Figure 9: (Color Online) Energy Spectra of each crystal before and after reconstruction.
In order to understand the penetrating cosmic-ray muons, which deposit large amounts of energy in the detectors and mix with the high energy $\gamma$, we examined the spatial distribution of signals in each event. The spatial spreading of the signals in the horizontal ($x$) and vertical ($y$) directions is defined as follows:
$\delta x$ $=\sum E_{i}|x_{i}-\bar{x}|/E_{\text{tot}},$ (6)
$\delta y$ $=\sum E_{i}|y_{i}-\bar{y}|/E_{\text{tot}}.$
Where $E_{i}$, $x_{i}$, and $y_{i}$ represent the energy, horizontal position, and vertical position of the $i^{\rm th}$ fired crystal unit, respectively. $E_{\text{tot}}$ denotes the total deposited energy, and $\bar{x}$ and $\bar{y}$ are the energy-weighted centroids of the incident $\gamma$-rays in the horizontal and vertical directions, respectively. Clearly, when only a single crystal unit is fired, both $\delta_{x}$ and $\delta_{y}$ are zero. A value of $\delta_{x(y)}=3.5~\text{cm}$ typically corresponds to events where two adjacent detectors are fired with nearly equal energy deposition.
Fig. 10 shows the scatter plots of $\delta_{y}$ versus $\delta_{x}$ for events in two energy intervals: (a) $10\leq E_{\text{tot}}\leq 100~\text{MeV}$, and (b) $E_{\text{tot}}>100~\text{MeV}$. Here the veto of the surrounding scintillators is already included in order to display the feature of the reconstructed events (same for the following figures 11 to 15). Even though the veto detectors are at work, there is still cosmic ray muons being recorded. The two distributions exhibit distinct features. For lower-energy events with $10\leq E_{\text{tot}}\leq 100~\text{MeV}$ (panel a), the distribution is nearly symmetric in the $x$ and $y$ directions, while for higher-energy events $E_{\text{tot}}>100~\text{MeV}$(panel b), the distribution is clearly dominated by events with large $\delta_{y}$, consistent with the picture that the cosmic-ray muons penetrate the hodoscope from above.
[…]
Fig. 11 further illustrates the correlation between the total energy $E_{\text{tot}}$ and the vertical spatial spread $\delta_{y}$ (panel a), as well as the correlation between $E_{\text{tot}}$ and the overall spatial spread $\delta r=\sqrt{\delta_{x}^{2}+\delta_{y}^{2}}$ (panel b). The events are clearly separated into two distinct groups. The high-energy group corresponds to cosmic-ray muons that are randomly coincident with the reaction events within the trigger window. In contrast, the low-$E_{\text{tot}}$ group is dominated by $\gamma$-rays originating from HICs. The $\gamma$-ray signals from the reactions are primarily deposited in one or two scintillator units, showing spread features of either $\delta_{y}=0$ or $0<\delta_{y}<3.5~\text{cm}$ ($\delta r=0$ or $0<\delta r<5~\text{cm}$), consistent with localized energy deposition. Some events with total energy below 20 MeV involve more than two fired units, likely originating from multiple statistical $\gamma$ emissions.
Figure 11: (Color Online) Correlations between the total energy $E_{\text{tot}}$ and the spatial spread: (a) $\delta_{y}$ versus $E_{\text{tot}}$ (b) $\delta r=\sqrt{\delta_{x}^{2}+\delta_{y}^{2}}$ versus $E_{\text{tot}}$.
The way that the total energy distribute in the fired CsI(Tl) units differs in $\gamma$-ray events and cosmic-ray muon events. According to the development procedure of an electromagnetic shower, most of the energy are deposited in the event core of the high energy incident $\gamma$. For the cosmic-ray muon event, on the other hand, each fired CsI(Tl) unit on the path tends to record similar energy deposit. Fig. 12 presents the correlation between the energy deposit of the event core $E_{\rm core}$ and $E_{\rm tot}$ to distinguish the two types of events. Panels (a) and (b) correspond to events fired by central units and edge units, respectively. It can be seen that high energy $\gamma$ events and the cosmic ray muon events can be separated. For the cosmic ray muons situating at right half of the plot, the $E_{\rm tot}$ is much higher than $E_{\rm core}$, indicating there are several units containing similar large energy deposit as the core unit. For the $\gamma$ events within the primary energy range of bremsstrahlung $\gamma$-rays namely from tens MeV to approximately 100 MeV, the feature of $E_{\rm tot}\gtrapprox E_{\rm core}$ is visible, indicating a significant portion of the total energy is deposited in the central fired unit, with only a small fraction leaking into neighboring detectors. At the low energy end below 20 MeV, $E_{\rm tot}$ is again far towards right side of $E_{\rm tot}=E_{\rm core}$, in accordance with the speculation that these events correspond to the emissions of multiple statistic $\gamma$-rays at low energies, which are not of interest of current analysis.
Figure 12: (Color Online) Correlations between $E_{\rm core}$ and $E_{\rm tot}$ of the events. (a) Events with central units as the fire center. (b) Events with out ring as the fire center.
III.5 Decomposition of Trigger Conditions
The experimental data acquisition is triggered by various conditions as described in section II. The energies recorded in the CsI(Tl) hodoscope may or may not be counted in the trigger conditions. It is therefore important to check the $\gamma$ spectrum at different trigger conditions.
We start with the correlation between the total energy and the timing information extracted from the core unit of the CsI(Tl) hodoscope, as shown in Fig. 13. The abscissa is the time information of the event core after doing the A-T correction, while the ordinate represents $E_{\text{tot}}$. Some interesting features are visible from the figure. 1) The amplitude dependence of the timing signal has been effectively removed, particularly at low energies. 2) The dense region covers the time range between $300<T_{\gamma}<-150~\text{ns}$. And within this range, two straight bands are clearly visible for the high energy events. These two bands have clear origin (see next). 3) A noticeable distribution appears outside the main time window. The broad distribution of $T_{\gamma}$ can not be attibuted to the slow time response of the CsI crystals only. To gain a clearer understanding of the energy-time (E-T) correlation characteristics, it is therefore essential to examine the E-T distributions separately for each trigger type.
Figure 13: (Color Online) Correlations between the reconstructed $\gamma$ energy and the timing information of the core CsI(Tl) crystal unit.
In our experiment, each individual trigger signal is fed to TDC after certain delays. Once a trigger signal is generated, this trigger signal under inspection appears at the corresponding self-trigger peak position on its TDC channel distribution. If the trigger is generated by other conditions, the signal under inspection deviates from its self-trigger peak. Fig.14 shows the Trigger TDC distributions for various trigger modes: (a) SSD M1 & CsI M1, (b) SSD M2, (c) SSD M1 & NA M1, (d) NA M1 & T0, (e) LS & T0, and (f) ALL_OR (global trigger). In the TDC distribution of the global trigger in panel (f), a series of small periodic peaks can be observed in the later time region, which originate from trigger signals in different high frequency cycles, with a period of 74 ns. For each individual trigger mode, as shown in panels (a) to (e), a sharp and prominent main peak is presented on the corresponding TDC spectrum. By selecting events within this main peak, one can obtain a clean sample of events triggered exclusively by the corresponding mode. Due to the limited resources of available TDC readout channels, the TDC signal for the SSD M1 & LS  trigger was not recorded during the experiment, but the missing of this specific channel does not influence our discussions.
Figure 14: (Color Online) TDC channel distributions for the trigger conditions. (a) SSD M1 & CsI M1, (b) SSD M2, (c) SSD M1 & NA M1, (d) NA M1 & T0, (e) LS & T0, and (f) ALL_OR (global trigger).
The correlations between the reconstructed energy $E_{\rm tot}$ from the CSHINE-Gamma detector and $T_{\text{core}}$ under different trigger conditions are shown in Fig. 15. Panels (a)–(e) correspond to five distinct trigger modes same as in Fig. 14. In panel (a), the trigger condition is set by requiring both a track found in SSDT and a hit in the $\gamma$ hodoscope, with the latter being designed to determine the trigger timing peak. Clearly seen, the self-trigger peak is clearly visible at the left timing. Besides, there are also dense scattering points in the main coincidence window, indicating that trigger timing peak is determined by the SSD M1  signal, which exhibits large time jitter because the SSDT signal is generated from the shaping amplifiers.
From Fig. 15 (b) corresponding to SSD M2, it is shown that the scattering points for the high energy $\gamma$-rays are populated broadly in the entire TDC range, suggesting that the broad time distribution discussed in Fig. 13 is primarily contributed by the SSD M2  trigger mode. The diffuse distribution over time does not mean the high possibility of random coincidence, nor the poor timing resolution of the CsI(Tl) detectors for the high energy $\gamma$-rays. Otherwise because similar feature shall appear in other trigger conditions. Instead, as illustrated in Fig. 2, in this mode, the MULT analog signal from the shaping amplifiers of all the SSDTs are added by an analog fan-in/fan-out module (N625) before being sent to CF8000 module, which further delivers a logical NIM signal as the trigger condition SSD M2  by setting a discrimination threshold of 70 mV (equivalent to multiplicity $M\geq 2$). Since the input signal fed to the CF8000 is a step-like trigger signals with very uncertain leading edge, the trigger timing window suffers large jitter, and thus the timing signal of the CsI(Tl) recorded in TDC exhibits large variation too. The same reason can be applied to explain the diffuse distribution in panel (a) in the main coincidence window. It suggests that if one includes this part of data in the analysis of the bremsstrahlung $\gamma$-rays, it is required to use the slow coincidence window, instead of the fast one.
Fig. 15 (d) and (e) further confirm the speculation mentioned above. In both trigger conditions of NA M1 & T0  and LS & T0, the $\rm BaF_{2}~T_{0}$ detectors are included in trigger calculation and determines the trigger timing, since both the CsI(Tl) and the $\rm BaF_{2}$ response to $\gamma$-rays (here we neglect to mention the mouns without influencing the discussion), the timing signal of event core $T_{\rm core}$ keeps nearly constant, as clearly shown by the vertical band at large $E_{\rm core}$. The cleanness of the high energy events in panel (c) and (d) is consistent with the speculation that the scattering distribution in panel (b) is mainly attributed to the poor timing of the trigger signal of SSD M2. In panel (c) corresponding to SSD M1 & NA M1 , there are fewer high energy $\gamma$ events.
Last, we can look back the two vertical bands observed in Fig. 13. The left band in Fig. 13 originates entirely from events triggered by the self-trigger mode of SSD M1 & CsI M1, where $\gamma$ hodoscope determines the trigger timing. In contrast, the right band corresponds to events triggered by the conditions involving the $\rm T_{0}$ detector. Since both bands represent $\gamma$ events with constant velocity, the time difference between them is attributed to electronics settings, since both detectors recorded $\gamma$-rays.
Figure 15: (Color Online) Correlations between reconstructed energy $E_{\rm tot}$ from CSHINE-Gamma and the timing information of the core CsI(Tl) crystal with the different trigger conditions.
III.6 Detector Response Matrix
The detector response matrix is constructed using Geant4 simulations, where each element $P_{i}(E_{j})$ denotes the probability of an incident photon with energy $E_{j}$ depositing an energy $E_{i}$ which is reconstructed in the $\gamma$ hodoscope. The matrix spans photon energies from 1 MeV to 130 MeV, with 1 MeV binning, resulting in a $130\times 130$ matrix. The whole matrix is normalized to ensure proper probabilistic interpretation. For visualization, Fig. 16 shows the distribution of matrix elements $P(E_{\text{ini}},E_{\text{out}})$, using only the leading $100\times 100$ subset. This response matrix plays two essential roles in the analysis: (1) forward-folding theoretical energy spectra for comparison with experimental data, and (2) unfolding measured spectra to recover the original photon energy distribution.
Figure 16: (Color Online) The distribution of the detector response matrix elements $P(E_{\rm ini},E_{\rm out})$. Here $P(E_{\rm ini},E_{\rm out})$ is the possibility of recognizing an incident $\gamma$-ray with initial $E_{\rm ini}$ to be $E_{\rm out}$ as the detector output.
III.7 Cosmic-ray Muon Background
Although three plastic scintillators are mounted surrounding the $\gamma$ hodoscope, the high energy $\gamma$ spectrum is inevitably contaminated by cosmic ray muons. To subtract the contamination, it is required to measure the cosmic ray muons. During beam-off periods, a series of background $\gamma$-ray events were recorded. These events are primarily composed of cosmic-ray muons detected by the CSHINE-Gamma array and low-energy residual $\gamma$-rays. By normalizing the beam-off spectra to the total number of cosmic-ray events after events reconstruction, one can subtract the beam-off background from the beam-on measured spectra. This procedure yields a clean $\gamma$-ray energy spectrum originating exclusively from the HICs.
To gain the first view differentiating the cosmic ray muon and the high energy $\gamma$-ray, Fig. 17 presents two representative event displays. Panel (a) shows a typical high-energy $\gamma$-ray event, where the majority of the energy is deposited in a single scintillator module, with a small amount of leakage into adjacent modules. Panel (b) displays a typical cosmic muon event. The muon traverses the entire CSHINE-Gamma array from top to bottom, depositing energy in each scintillator along its path. The amount of energy deposited in each unit correlates with the muon’s path length through that unit, allowing a clear visualization of the muon’s trajectory across the array.
Figure 17: (Color Online) Event display for (a) a typical $\gamma$ event and (b) a typical cosmic ray event.
We now examine in detail the characteristics of events recorded during the beam-off background measurements. The correlation between $\delta y$ and $\delta x$ is shown in Fig. 18(a). Most events are concentrated in regions with large $\delta y$, consistent with the vertical trajectories expected from cosmic-ray muons. The correlation between the core crystal energy ($E_{\rm core}$) and the total deposited energy ($E_{\rm tot}$) is shown in Fig. 18 (b). In the higher $E_{\rm tot}$ range, a significant number of events exhibit relatively low energy deposition in the event core scintillator, indicating that the energy is more evenly distributed among multiple crystals, again aligned with the characteristics of through-going cosmic-ray muons.
Figure 18: (Color Online) (a) Correlations between $\delta y$ and $\delta x$ of the cosmic-ray muon events. (b) Correlations between $E_{\rm core}$ and $E_{\rm tot}$ of the cosmic-ray muon events.
IV IBUU-MDI Model Description
IV.1 Model Description
The Isospin-dependent Boltzmann-Uehling-Uhlenbeck (IBUU) transport model [74, 75, 76, 56, 77] is employed in the theoretical calculations, which effectively describes the dynamics of nucleon-nucleon (NN) collisions, with its main equation given by
$\frac{\partial f}{\partial t}+\vec{v}\cdot\nabla_{r}f-\nabla_{r}U\cdot\nabla_{p}f=I_{\rm coll},$ (7)
where $f(\vec{r},\vec{p},t)$ is the probability of finding a particle at time $t$, with momentum $\vec{p}$ at position $\vec{r}$. $U$ represents the mean-field potential, and the evolution of $f(\vec{r},\vec{p},t)$ in elastic and inelastic two-body collisions is governed by the collision term $I_{\rm{coll}}$,
$I_{\rm{coll}}=$ $-\frac{1}{(2\pi)^{3}}\int\mathrm{d}^{3}p_{2}\mathrm{d}^{3}p_{2^{{}^{\prime}}}\mathrm{d}\Omega\frac{\mathrm{d}\sigma}{\mathrm{d}\Omega}v_{12}$ (8)
$\times[ff_{2}(1-f_{1^{{}^{\prime}}})(1-f_{2^{{}^{\prime}}})-f_{1^{{}^{\prime}}}f_{2^{{}^{\prime}}}(1-f)(1-f_{2})]$
$\times\delta^{3}(\vec{p}_{1}+\vec{p}_{2}-\vec{p}_{1^{{}^{\prime}}}-\vec{p}_{2^{{}^{\prime}}}),$
where $\frac{\rm{d}\sigma}{\rm{d}\Omega}$ is the in-medium NN cross section and $v_{12}$ is the relative velocity of the two colliding nucleons. The coordinates $(\vec{r}_{1},\vec{p}_{1})$ and $(\vec{r}_{2},\vec{p}_{2})$ refer to the phase-space positions of nucleon 1 and nucleon 2 before collision, and change to $(\vec{r}_{1},\vec{p}_{1^{{}^{\prime}}})$ and $(\vec{r}_{2},\vec{p}_{2^{{}^{\prime}}})$ after collision. The Pauli blocking effect is also taken into account. The scattering is allowed only if the phase-spaces around $(\vec{r}_{1},\vec{p}_{1^{{}^{\prime}}})$ and $(\vec{r}_{2},\vec{p}_{2^{{}^{\prime}}})$ are unoccupied; otherwise, the scattering is suppressed [77]. This effect is embodied in the last term of $I_{\rm{coll}}$.
Moreover, one of the most important inputs in the IBUU model is the the mean-field potential. We adopt in this work the momentum-dependent interaction (MDI) potential derived from the Gogny effective interaction [57],
$U(\rho,\delta,\vec{p},\tau)=$ $A_{u}(x)\frac{\rho_{\tau^{\prime}}}{\rho_{0}}+A_{l}(x)\frac{\rho_{\tau}}{\rho_{0}}$ (9)
$+B(\frac{\rho}{\rho_{0}})^{\sigma}(1-x\delta^{2})-8x\tau\frac{B}{\sigma+1}\frac{\rho^{\sigma-1}}{\rho^{\sigma}_{0}}\delta\rho_{\tau^{\prime}}$
$+\frac{2C_{\tau,\tau}}{\rho_{0}}\int\mathrm{d}^{3}\vec{p^{\prime}}\frac{f_{\tau}(\vec{r},\vec{p^{\prime}})}{1+(\vec{p}-\vec{p^{\prime}})^{2}/\Lambda^{2}}$
$+\frac{2C_{\tau,\tau^{\prime}}}{\rho_{0}}\int\mathrm{d}^{3}\vec{p^{\prime}}\frac{f_{\tau^{\prime}}(\vec{r},\vec{p^{\prime}})}{1+(\vec{p}-\vec{p^{\prime}})^{2}/\Lambda^{2}},$
where $\tau=\pm 1/2$ denotes the isospin (+1/2+1/2 for neutron and 1/2-1/2 for proton). $\rho_{n}$ and $\rho_{p}$ are the neutron and proton densities. $\rho=\rho_{n}+\rho_{p}$ is the total nucleon density. $\rho_{0}=0.16\,\rm{fm}^{-3}$ is the saturation density, and $\delta=(\rho_{n}-\rho_{p})/(\rho_{n}+\rho_{p})$ is the isospin asymmetry. The parameters $A_{u}(x)$, $A_{l}(x)$, $B$, $C_{\tau,\tau}$, $C_{\tau,\tau^{\prime}}$, $\sigma$ and $\Lambda$ are taken from Ref.[57]. Notably, the different choices of $x$ parameter correspond to different density dependence of the symmetry energy.
For comparison, a momentum-independent soft Bertsch-Kruse-Das Gupta (SBKD) potential is also used in the IBUU simulations [78],
$U(\rho)=A(\frac{\rho}{\rho_{0}})+B(\frac{\rho}{\rho_{0}})^{\sigma},$ (10)
where $A$, $B$ and $\sigma$ can be expressed in terms of the nuclear incompressibility coefficient in the equation of state [79].
Another important input in the transport model is the NN elastic cross section. In IBUU, the isospin-dependent in-medium NN cross section using nucleon effective mass is given by [80]
$\sigma^{\rm medium}_{\rm NN}=\sigma^{\rm free}_{\rm NN}(\frac{\mu_{\rm NN}^{\ast}}{\mu_{\rm NN}})^{2},$ (11)
where $\mu_{\rm NN}$ and $\mu_{\rm NN}^{\ast}$ are the free-space reduced mass and in-medium reduced mass of the colliding nucleon pair, respectively. $\sigma^{\rm free}_{\rm NN}$ is the free-space NN cross section taken from experimental data [81].
The IBUU simulations utilize the initial single-nucleon momentum distribution $n(k)$ that incorporates the SRC and HMT effects [56, 53, 82]. For the symmetric nuclear matter (SNM), the HMT induced by SRC [83] is given by
$n_{\rm SNM}(k)=\left\{\begin{array}[]{rcl}A,&&{k\leq k_{\rm F}}\\ C/k^{4},&&{k_{\rm F}<k\leq\lambda k_{\rm F}}\\ 0,&&{k>\lambda k_{\rm F}},\end{array}\right$. (12)
where $k$ represents the single nucleon momentum. $k_{\rm F}$ is Fermi momentum and $\lambda$ is the high-momentum cutoff [53, 82]. The parameters $A$ and $C$ are determined by the normalization condition,
[…]
For the asymmetric nuclear matter (ANM), the high-momentum component is linearly dependent on the isospin asymmetry [84, 85, 86]. The $n(k)$ for ANM is parameterized as
$n_{\rm ANM}(k)_{p}=\left\{\begin{array}[]{rcl}B,&&{k\leq k^{p}_{\rm F}}\\ C(1+\delta)/k^{4},&&{k^{p}_{\rm F}<k\leq\lambda k^{p}_{\rm F}}\\ 0,&&{k>\lambda k^{p}_{\rm F}},\end{array}\right$. (14)
$n_{\rm ANM}(k){{}_{n}}=\left\{\begin{array}[]{rcl}B{{}^{\prime}},&&{k\leq k^{n}_{\rm F}}\\ C(1-\delta)/k^{4},&&{k^{n}_{\rm F}<k\leq\lambda k^{n}_{\rm F}}\\ 0,&&{k>\lambda k^{n}_{\rm F}},\end{array}\right$. (15)
where the parameters $B$ and $B{{}^{\prime}}$ are determined by the normalization condition,
$4\pi\int\limits_{0}^{\infty}n_{\rm ANM}(k)_{p(n)}k^{2}dk=1.$ (16)
The bremsstrahlung reaction channel is incorporated into the collision processes within the IBUU model, which employs the Bertsch criterion [87] and Pauli exclusion principle to determine whether a NN collision occurs. Bremsstrahlung photons are generated through two channels, the neutron-proton collisions ($np\gamma$) and the proton-proton collisions ($pp\gamma$). However, the $np\gamma$ channel dominates as the dipole moment in a charge-symmetric $pp$ system vanishes. Consequently, the lowest-order multipole radiation in the $pp$ collisions is of quadrupole type, whose intensity is much lower compared to the dipole radiation in the $np$ collisions [88, 89]. Hence we consider only the $np\gamma$ process in our calculations.
The bremsstrahlung photons emitted in the $np$ collisions are produced in the processes of deceleration and acceleration of the protons interacting with the neutrons. In the intermediate energy range, we adopt the neutral scalar $\sigma$ meson exchange model [88] in the present work, in which the photons are emitted via the external processes, illustrated by the Feynman diagrams in Fig.19 (a). At higher energies, one will have to take the charged meson exchange processes into consideration. The photons can also be emitted from the internal mesons, to which the electromagnetic field also couples [88, 90], and the Feynman diagram is given by Fig.19 (b).
Figure 19: Feynman diagrams of bremsstrahlung photons in npnp scattering from (a) external lines and (b) internal lines.
Due to the low probability of producing bremsstrahlung photons in the reaction, the impact of bremsstrahlung on nucleon kinematics is also negligible, and hence, the perturbation method is applied to calculate the probability of photon production in each $np$ collision. Based on the $\sigma$ meson exchange model mentioned above, a well fitted expression for the probability of elementary double differential photon production can be applied in IBUU simulations [91],
$\frac{{\rm d^{2}}P}{{\rm d}\Omega{\rm d}E_{\gamma}}=1.671\times 10^{-7}\frac{[1-(E_{\gamma}/E_{\rm max})^{2}]^{\alpha}}{E_{\gamma}/E_{\rm max}},$ (17)
where $E_{\gamma}$ represents the energy of produced photons, and $E_{\rm max}$ represents the total available energy in the center of mass frame of the colliding neutron-proton pair. The coefficient $\alpha=0.7319-0.5898\beta$, where $\beta$ represents the velocity of nucleon. The total photon production probability per event is the sum of the probabilities of all $np$ collisions producing photons throughout the entire process.
IV.2 Parameter influence test
In the IBUU simulations, one has to choose several important inputs, which may influence the results of final observables. Here we systematically check the sensitivity of bremsstrahlung photon production on the choices of parameters, namely the impact parameter, the nuclear mean-field potential, and the symmetry energy parameter.
Firstly, we show in Fig. 20 the variations of bremsstrahlung photon production by using different impact parameters. We adopt $b=0\sim 5\,\rm{fm}$ in our IBUU simulations. To check the sensitivity of photon production on $b$ parameter, several other values are used, namely, $b=0\sim 9\,\rm{fm}$ and $b=0,5,9\,\rm{fm}$. Their corresponding spectra are normalized to the $b=0\sim 5\,\rm{fm}$ one. This is because we focus on the spectral shape similarities rather than absolute normalization. Although the absolute values of bremsstrahlung photon production spectra vary with different choices of impact parameters, but the spectral shape remain almost the same to the impact parameters, as seen in Fig. 20.
Figure 20: The double differential photon production calculated by varying impact parameters. The impact parameters are taken to be $b=0\sim 5\,\rm{fm}$, $b=0\sim 9\,\rm{fm}$ and fixed values $b=0,5,9\,\rm{fm}$. The results using $b=0\sim 9,0,5,9\,\rm{fm}$ are normalized to the $b=0\sim 5\,\rm{fm}$ case.
In Fig. 21, the variations of bremsstrahlung photon production by using different nuclear mean-field potentials are shown. We adopt the MDI potential with the symmetry energy parameter x=1x=-1 [66]. The momentum-independent SBKD potential is used to check the sensitivity of mean-field potentials, and the parameters used in SBKD potential are listed in Table3. The $\sigma$ parameter represents the softness or stiffness of the potential with the variation of the nuclear density [78]. For the possible influence of symmetry energy parameters in MDI, we adopt three different parameters x=1x=-1, $x=0$ and $x=1$ in simulations (see Fig. 22). The tests have shown that within the energy region investigated in this work, our conclusions are not affected by the choice of nuclear potentials or the symmetry energy parameters. This is probably because the bremsstrahlung photons interact with nucleons only electromagnetically, and are not sensitive to the isoscalar and isovector parts of nuclear interactions.
Figure 21: The double differential photon production calculated by varying the choice of nuclear mean-field potentials. We compare the results from the MDI potential with those from the momentum-independent SBKD potential. The results from the SBKD potential are normalized to the MDI results.
Table 3: The parameters $\sigma$, $A\,{\rm{(MeV)}}$ and $B\,{\rm{(MeV)}}$ used in the momentum-independent SBKD potential.
$\sigma$ $A\,{\rm{(MeV)}}$ $B\,{\rm{(MeV)}}$
7/6 -356 303
4/3 -218 164
2 -124 70.5
2.77 -103.22 49.56
Figure 22: The double differential photon production calculated by varying symmetry energy parameters. Three different $x$ parameters are employed in the simulations, namely x=1x=-1, $x=0$ and $x=1$. The results using $x=0$ and $x=1$ are normalized to the x=1x=-1 case.
IV.3 Reference Frame Transformation
The mass of each nulceon is $m_{\rm u}=931.5~\rm MeV$ and the kinetic energy of each nulceon is $E_{\rm u}=25~\rm MeV$. The momentum of each nulceon in the beam is
$p_{\rm u}=\sqrt{(E_{\rm u}+m_{\rm u})^{2}-m_{\rm u}^{2}},$ (18)
the rapidity of the beam can be obtained by $p_{u}$ as the following,
$sinhy_{b}=\frac{p_{\rm u}}{m_{\rm u}}.$ (19)
In the system of 124Sn+124Sn, the mass of the nuclei in the beam $M_{\rm b}=124$ and the mass of the target nuclei $M_{\rm t}=124$ can be used for the calculation of the rapidity in the center-of-mass frame system,
$y_{\rm cm}=\frac{1}{2}y_{\rm b}+\frac{1}{2}\ln[\frac{M_{\rm b}\exp(y_{\rm b})+M_{\rm t}}{M_{\rm b}+\exp(y_{\rm b})M_{\rm t}}].$ (20)
And then, we can obtain the velocity in the center-of-mass frame frame,
$\beta_{\rm cm}=\frac{\exp(2y_{\rm cm})-1}{\exp(2y_{\rm cm})+1}.$ (21)
The $\gamma$ detector array is placed at $\theta_{\rm b}=110^{\circ}$ with respect to the beam direction, and the transformation factor can be finally obtained as the following,
$L(\theta_{\rm b})=\frac{1-\beta_{\rm cm}\cos\theta_{\rm b}}{\sqrt{1-\beta_{\rm cm}^{2}}}=1.04631.$ (22)
The photon Doppler effect causes softening of the energy spectrum. During the conversion of the spectrum from the laboratory frame to the center-of-mass frame, the energy must be multiplied by the factor $L$, while the counts are divided by the same factor. The experimentally detected $\gamma$-ray spectrum can be transformed into the center-of-mass frame using this method.
Since the detected $\gamma$-ray spectrum in the experiment is influenced by the detector filter, theoretical curves must also account for the detector filter when being compared with experimental data. Using Geant4 simulations, we modeled the CsI array’s response to $\gamma$-rays of varying energies and obtained the detector filter matrix in the laboratory frame, as depicted in Fig. 16. To incorporate this effect, the theoretical curves were first transformed into the laboratory frame, passed through the detector filter matrix, and then converted back to the center-of-mass frame. This procedure enables a direct comparison between the experimental data and the IBUU theoretical predictions in subsequent analyses.
V Results and Discussions
V.1 Analysis of Systematic Uncertainty
The systematic uncertainties primarily stem from three sources: (1) variations in the spectra obtained from radioactive source calibrations on different dates during the experiment, (2) the choice of normalization energy range for high-energy cosmic rays during background subtraction, and (3) the linear assumption of the detector response at higher $\gamma$ energy levels. Analysis shows that the choice of normalization energy ranges exhibit no significant variations, indicating that the effect of cosmic ray normalization range on the $\gamma$-ray spectrum is negligible.
The experimental data included three sets of linear calibration obtained from radioactive sources on different dates. Among these, the $\gamma$ energy spectrum calibrated using the dataset with the highest statistics was designated as the central spectrum. In addition to the linear calibration, we introduced two alternative detector response correction functions (DRCFs) characterizing the differences between the linear calibration and the actual quadratic calibration obtained on SLEGS test experiment [72]. Then a total of nine distinct $\gamma$ energy spectra can be derived. Excluding the central spectrum, we calculated the standard deviation of the remaining eight spectra for each bin as the systematic uncertainty. During the evaluation of the likelihood functions and $\chi^{2}$ (see next), the values derived from the central spectrum were used as the central points. The likelihood function and $\chi^{2}$ values computed from the remaining spectra were utilized to determine the associated errors (standard deviations). This methodology enabled us to obtain likelihood and $\chi^{2}$ functions with systematic uncertainties at different $R_{\text{HMT}}$ values.
V.2 Experimental $\gamma$ spectrum in slow coincidence
As shown by the primary analysis, the method introduced in the previous experiment [66] is applied. All the $\gamma$ events in the slow coincidence window are counted. Namely a valid timing of the reconstructed $\gamma$ is found within the TDC range. The background is obtained by the beam-off measurement. The experimental data calibration and background subtraction were performed using the background data and radioactive source calibration data on March 8, 2024. Calibration files were generated based on this data and subsequently applied to all experimental and background data for calibration and reconstruction. This process yielded the total beam-on spectrum and the beam-of background spectrum, shown as the black and red lines in Fig. 23(a), respectively. The red background spectrum was scaled by a normalization factor $r_{\rm n}$. Specifically, the total counts above $110\rm MeV$ in both the total spectrum ($Y_{\rm on}$) and the background spectrum ($Y_{\rm off}$) were calculated, and the yield ratio is defined as $r_{\rm n}=\frac{Y_{\rm on}(E_{1})}{Y_{\rm off}(E_{1})}$, where the subscript ‘on’ and ‘off’ represent the beam-on and beam-off spectrum, respectively. $E_{1}=110$ MeV is the low energy border taken to do the normalization, the upper border is set to 200 MeV. Then the ratio $r_{\rm n}$ is applied to the entire background spectrum, resulting in the red histogram in Fig. 23(a). By subtracting the scaled background spectrum from the total energy spectrum, the background-subtracted $\gamma$ spectrum was obtained, as shown in Fig. 23(b), where we take 01000-100 MeV as the energy range of total $\gamma$-rays.
Figure 23: (Color Online) (a) Total energy Spectrum of beam-on (black) in all trigger conditions and beam-off measurement (red). (b) The $\gamma$ energy spectrum after subtracting the background.
[…]
Figure 24: (Color Online) Comparison of reconstructed energy spectra. (a) Energy spectra reconstructed from the central four units, edge six units, and all units combined. (b) The same spectra after normalization in the $35\sim 100$ MeV energy range.
Fig. 25 illustrates the rebinned experimental $\gamma$ spectrum in the c.m. frame. The black dots represent the central spectrum, with statistical uncertainties indicated by error bars, and the gray shaded regions denote the systematic uncertainties at each energy point. Several key theoretical curves, processed with the detector filter, are presented for comparison. Unlike in our previous measurement [66, 65], in this experiment, the $\gamma$ hodoscope was operated in active triggering mode, preventing the determination of the number of $np$ collisions in heavy-ion reactions as a normalization factor for the $\gamma$ spectrum. Consequently, the comparison between experimental data and theoretical curves predicted by the IBUU model focused on spectral shape similarities rather than absolute normalization. To align the range of the theoretical curves with the experimental spectrum, the theoretical values were uniformly scaled by a factor of $2.5\times 10^{9}$, which had no impact on determining the $R_{\rm HMT}$, as it did not affect the comparison of shape differences between the experimental and model-predicted spectra.
Figure 25: (Color Online) Comparison of the rebinned experimental $\gamma$ spectrum (black dots) with statistical uncertainties (error bars) and systematic uncertainties (gray shaded areas) in the c.m. frame. Several key theoretical curves, processed with the detector filter, are overlaid for comparison.
Energies below $30~\rm MeV$ are primarily influenced by collective resonance and statistical emissions as well as the random coincidence. Therefore, we selected an energy range of $35<E_{\gamma}^{\rm cm}<100~\rm MeV$ as the central analysis interval and calculated the likelihood function between the experimental $\gamma$ spectrum at each point within this range and various IBUU theoretical curves.
The comparison between the IBUU model and the experimental data was performed using the maximum likelihood analysis method, which assumes that the total counts are fixed and that the experimental spectrum can be treated as a histogram, with each bin following a multinomial distribution. The likelihood function can be defined as follows, primarily focusing on the similarity in the shape of the spectra,
$L(R_{\text{HMT}})$ $=n!\prod_{i}^{\text{range}}\frac{1}{n_{i}!}p_{i}^{n_{i}}(R_{\text{HMT}}),$ (23)
$\ln L(R_{\text{HMT}})$ $=\sum_{i}^{\text{range}}n_{i}\ln p_{i}(R_{\text{HMT}})-\sum_{i}^{\text{range}}\ln n_{i}!+\ln n!,$
where $i$ represents for the sum of experimental points within a certain statistical analysis interval and $n_{i}$ is the counts of $i^{\rm th}$ experimental data point. $p_{i}$ represents the probability that the theoretical model predictions falling within the corresponding histogram bin of the experimental data under the specified statistical analysis interval and for a given $R_{\text{HMT}}$. Additionally, in the given statistical interval, all the $p_{i}$ should be normalized. The second and third terms are model independent, which can be neglected to simplify the calculation. Thus, a simplified logarithmic likelihood function can be defined as
$\ln L^{\prime}(R_{\text{HMT}})=\sum_{i}^{\text{range}}n_{i}\ln p_{i}(R_{\text{HMT}}).$ (24)
The results are presented as red dots with error bars in Fig. 26, where the central values are derived from the central $\gamma$ spectrum. To standardize the likelihood function values obtained from different $\gamma$ spectra, the likelihood function value corresponding to the theoretical curve derived from $R_{\text{HMT}}=0\%$ was selected as the reference point ($\ln L(R_{\text{HMT}}=0\%)=0$). The relative likelihood function values, $\Delta\ln L^{\prime}(R_{\text{HMT}})$, for all other points were calculated relative to this baseline in each spectrum. The error bars at each $R_{\text{HMT}}$ represent the standard deviations of the relative likelihood values obtained from various $\gamma$ spectra generated using different calibration parameters in comparison to the model predictions with different $R_{\rm HMT}$.
The likelihood function reaches its maximum near $R_{\text{HMT}}=20\%$ and follows an approximately quadratic distribution. The trend was fitted with a quadratic function within the range of $5\%<R_{\text{HMT}}<35\%$, depicted as the dashed curve in the figure, which also yielded a maximum value at $R_{\text{HMT}}=(20.1\pm 0.4)\%$, where the uncertainty of 0.4% reflects the fitting error from the quadratic model.
Figure 26: The likelyhood function values of different $R_{\text{HMT}}$ and the corresponding quadratic fitting in the energy range of $35~\rm MeV$ to $100~\rm MeV$.
To determine the confidence intervals for the HMT ratio based on the likelihood function distribution, $R_{\text{HMT}}$ values were identified that correspond to relative likelihood values smaller than the maximum by $\frac{m^{2}}{2}$, which represent confidence levels of $\pm m\sigma$. The $R_{\text{HMT}}$ ranges for likelihood values reduced by $0.5$, $4.5$ and $12.5$ from the quadratic maximum correspond to $\pm 1\sigma$, $\pm 3\sigma$ and $\pm 5\sigma$ confidence intervals, respectively, as indicated by the black dashed lines in the figure. Based on the experimentally measured $\gamma$ spectrum, the value of $R_{\text{HMT}}$ is determined at $1\sigma$ confidence level as
$R_{\text{HMT}}=(20.1\pm 2.1)\%.$ (25)
Worth mentioning, the spectrum obtained here does not show a broad hump-like structure near $E_{\gamma}\approx 60$ MeV, at variance with the $\gamma$ spectrum obtained in the previous Kr+Sn experiment at the same beam energy [66], where the statistics was much lower. According to the new result, the presence of $np\rightarrow d\gamma$ channel in the collision is not apparently supported.
V.3 Experimental $\gamma$ spectrum in fast coincidence
Alternatively, one can analyze the $\gamma$ spectrum in a different approach as a cross check, where both the $\gamma$ spectrum and the background are extracted in beam-on experiment. Specifically, the $\gamma$ events are accumulated in fast coincidence gate defined by the trigger, while the background is obtained in a time window with the same width but far away from the main trigger time window. By this approach, one ensures that only random coincidence are contained in the time window far away from the main trigger gate.
Noticeably, however, the events triggered by the SSD M2  condition shall be excluded now, because the $\gamma$ time are spreading across the entire TDC range due to serious timing jitter of the trigger signal. Otherwise, the truly coincident $\gamma$ events are recorded in the background window and the subtraction leads to a wrong energy spectrum.
To isolate true $\gamma$ events, we exclude all events of the SSD M2  trigger within the main peak region and define two time windows based on Fig. 13:
  • The events with $T_{\rm core}$ in the range $350~\text{ns}<T_{\rm core}<-50~\text{ns}$ are considered as true $\gamma$ events.
  • The events in the range $50~\text{ns}<T_{\rm core}<350~\text{ns}$ are taken as background.
The energy spectra constructed from these two time windows with equal width are shown in Fig. 27(a). The black curve represents the $\gamma$-ray energy spectrum within the fast coincidence window, while the red curve corresponds to the background window. Panel (b) shows the resulting $\gamma$ energy spectrum after subtracting the background directly. Since the time width is the same, no more normalization is needed.
Figure 27: (Color Online) (a) Total energy Spectrum in the fast coincidence window excluding SSD M2  trigger condition (black) and in the accidental coincidence window (red). (b) The $\gamma$ energy spectrum after subtracting the background.
The same comparison between the background-subtracted $\gamma$ spectrum (excluding the SSD M2  trigger) and theoretical model predictions, scaled by a factor of $1.7\times 10^{9}$, is shown in Fig. 28. Using a likelihood-based method, we compare the experimental spectrum with IBUU model predictions and extract an optimal $R_{\text{HMT}}=(18.0\pm 2.8)\%$. This value is consistent, within uncertainties, with the result obtained via beam-off background subtraction method, indicating that the extracted SRC fraction is robust and independent of the specific background subtraction method.
Figure 28: (Color Online) Comparison of the rebinned experimental $\gamma$ spectrum excluding SSD M2  trigger condition with key theoretical curves.
Due to the reduced statistics of photons, particularly above 60 MeV, even slight fluctuations in the high-energy region may influence the final result. Thus, the 2% deviation from the previous analysis is taken as a systematic uncertainty. Combined, the total uncertainty is $\sqrt{2.1^{2}+2^{2}+0.4^{2}}\approx 3$. As the final result, the fraction of HMT is written as
$R_{\text{HMT}}=(20\pm 3)\%.$ (26)
VI Reconstruction of the Original $\gamma$ Spectrum
Finally, in order to offer a $\gamma$ spectrum which can be compared to other theoretic calculations without knowing the response matrix, we also tried to reconstructed the original $\gamma$ spectrum from the measured one by solving an inverse problem. In addition, the results of $R_{\rm HMT}$ can also be checked one more time from the original spectrum. For this purpose, the Richardson Lucy (RL) algorithm, which has been widely applied in the nuclear physics [92, 93, 94, 95, 96, 97, 98], is adopted to solve the inverse problem [65].
VI.1 Richardson Lucy Algorithm
In optical deblurring problems, the true property $\mu$ of a photon is measured as $\nu$, with their distributions $\mathcal{F}(\mu)$ and $f(\nu)$ related by:
$f(\nu)=\int d\mu P(\nu|\mu)\mathcal{F}(\mu),$ (27)
where $P(\nu|\mu)$ represents the conditional probability of measuring a photon with real property $\mu$ as $\nu$.
Analogously, in HIC experiments, the true photon energy $\mu$ is detected as $\nu$ due to the detector filter. This allows the RL algorithm to be applied for reconstructing the original energy spectrum. In discretized form, the relationship between the distribution function of the real energy spectrum $\mathcal{E}(\mu)$ and the measured energy spectrum $e(\nu)$ can be related via the detector response matrix $D_{ij}$:
$e_{i}=\sum_{j}D_{ij}\mathcal{E}_{j},$ (28)
where $D_{ij}$ is obtained from Geant4 simulations with the given experimental setup and data analysis scheme. It quantifies the conditional probability that a photon with real energy $\mu$ will be detected as $\nu$.
The RL algorithm updates the estimated spectrum iteratively:
$\mathcal{E}_{i}^{(r+1)}=A_{i}^{(r)}\cdot\mathcal{E}_{i}^{(r)},$ (29)
with the amplification factor $A_{i}^{(r)}$
$A_{i}^{(r)}=\sum_{j}\frac{e_{j}}{e_{j}^{(r)}}T_{ji},$ (30)
where the predicted measurement at the $r$-th iteration $e^{(r)}$ is:
$e_{j}^{(r)}=\sum_{i}D_{ji}\mathcal{E}_{i}^{(r)}.$ (31)
The normalized transformation matrix $T_{ji}$ is defined as
$T_{ji}=\frac{W_{j}D_{ji}}{\sum_{j’}W_{j’}D_{j’i}},$ (32)
where $W_{j}=\sqrt{e_{j}}$ reflects Poisson statistics of the measured counts.
To assess statistical uncertainties, we perform multinomial sampling of the measured spectrum. Multiple pseudo-spectra are generated while preserving the total event count, and each is independently processed through the RL algorithm. The resulting ensemble of reconstructed spectra allows us to compute the standard deviation for each energy bin, which serves as our uncertainty estimate. In the RL reconstructed $\gamma$-ray spectrum, the central values are obtained by the measured spectrum, while the error bars reflect the propagated statistical uncertainties derived from the multinomial sampling. This method robustly accounts for statistical fluctuations in the measurement and their impact on the spectral reconstruction.
VI.2 Original Bremsstrahlung $\gamma$ Spectrum
The initial values for all iterative solutions of different spectra are set to the theoretical curve with $R_{\text{HMT}}=0\%$. This choice eliminates any prior bias toward a non-zero HMT ratio in the initial condition. By starting with $R_{\text{HMT}}=0\%$, the iterative process avoids the risk of a pre-imposed HMT contribution influencing the reconstructed spectrum. Furthermore, this initialization improves both the efficiency and accuracy of the iterations, ensuring that the final solution is derived purely from the input data without relying on an arbitrary starting condition.
The number of iterations is a critical factor in ensuring the convergence of the RL algorithm while maintaining computational efficiency. To determine the optimal stopping point, we monitor the standard deviation $\delta_{\text{exp}}$ between the experimental spectrum and the predicted spectrum (calculated using the detector response matrix) after each iteration. The quantity $\delta_{\text{exp}}$ is defined as:
$\delta_{\text{exp}}=\sum_{i}\left(\frac{e_{i}^{P}-e_{i}}{\sqrt{e_{i}}}\right)^{2}/N_{p},$ (33)
where $e_{i}$ represents the input measured value for the $i$th energy bin, $e_{i}^{P}$ is the predicted value from the current iteration, and $N_{p}$ is the total number of bins. Here, $N_{p}=100$, corresponding to the energy range of $1~\text{MeV}$ to $100~\text{MeV}$ with a step size of $1~\text{MeV}$.
The convergence condition is established by monitoring the relative change in $\delta_{\text{exp}}$ between consecutive iterations. The iteration is terminated when the rate of change of $\delta_{\text{exp}}$ becomes less than $1\times 10^{-2}$, as expressed by:
$\left|\frac{\delta_{\text{exp}}^{(r)}-\delta_{\text{exp}}^{(r-1)}}{\delta_{\text{exp}}^{(r-1)}}\right|<1\times 10^{-2}.$ (34)
This threshold is carefully chosen to strike a balance: avoiding premature termination caused by insufficient iterations and mitigating error amplification that may arise from excessive iterations.
The evolution of $\delta_{\text{exp}}$ as a function of iteration number in solving central spectrum is shown in Fig. 29(a). It can be observed that $\delta_{\text{exp}}$ gradually decreases with increasing iteration times, indicating a systematic improvement in the agreement between the experimental and predicted spectra. The end of curve is at the convergence point, signaling the optimal termination of the iterative process.
Figure 29: (a) The evolution of $\delta_{\text{exp}}$ as a function of iteration times in solving central spectrum. (b) The comparison between the reconstructed measured spectrum and the original measured spectrum in the experiment.
To verify the reliability of the reconstruction, the reconstructed spectrum in the lab frame is folded back using the detector filter matrix to obtain the corresponding predicted measured spectrum. This reconstructed measured spectrum is compared with the original central measured spectrum, as shown in Fig. 29(b). The close agreement between the two spectra demonstrates that the reconstructed real $\gamma$ spectrum is both accurate and efficient, validating the performance of the iterative algorithm.
After the iteration process described earlier, the reconstructed real $\gamma$ spectrum from the central measured spectrum is obtained and shown as black hollow dots in Fig. 30. The spectrum has been transformed from the lab frame to the center-of-mass frame and subsequently rebinned, combining every 4 bins into 1.
Figure 30: Comparison between several main theoretical curves and the reconstructed $\gamma$ spectrum.
With the reconstructed real $\gamma$ spectrum, we can directly compare the experimental results with IBUU theoretical predictions without requiring any additional corrections or assumptions.
To present the comparison more clearly, the main theoretical curves are scaled by a specific coefficient of $4.5\times 10^{9}$ and rebin number to match the experimental data format. The comparison between the reconstructed spectrum and the scaled theoretical curves is shown in Fig. 30.
This direct comparison allows us to quantitatively evaluate the theoretical predictions and determine the HMT ratio that best describes the experimental data.
VI.3 Direct Comparison to IBUU-MDI simulations
To determine the favored ratio of HMT from the reconstructed spectrum using the Richardson-Lucy (RL) method, we define the $\chi^{2}(R_{\text{HMT}})$ to quantify the agreement between the experimentally measured spectrum and the theoretical predictions:
$\chi^{2}(R_{\text{HMT}})=\sum_{i}\frac{1}{\sigma_{i}^{2}}(\text{counts}_{i}^{\text{exp}}-\text{counts}_{i}^{R_{\text{HMT}}})^{2},$ (35)
where $\text{counts}_{i}^{\text{exp}}$ and $\text{counts}_{i}^{R_{\text{HMT}}}$ represent the counts in the $i$th energy bin for the experimental data and the theoretical prediction with a given $R_{\text{HMT}}$, respectively, and $\sigma_{i}$ is the uncertainty of the reconstructed experimental spectrum in the $i^{\rm th}$ bin.
In addition, due to the need to compare the shape of the experimental data with the theoretical predictions, the theoretical curves must be scaled again when calculating the $\chi^{2}(R_{\text{HMT}})$. This scaling ensures that we minimize the value of $\chi^{2}(R_{\text{HMT}})$ for each theoretical curve, providing an accurate evaluation of the best match between the experimental and theoretical spectra. This procedure allows us to determine the optimal $R_{\text{HMT}}$ that best describes the data while accounting for differences in the overall amplitude between the experimental and theoretical curves. The theoretical curve that minimizes $\chi^{2}(R_{\text{HMT}})$ corresponds to the best-fit $R_{\text{HMT}}$, providing the most accurate description of the experimental data.
The energy range from $35~\rm MeV$ to $100~\rm MeV$ is chosen as the central analysis interval, and the $\chi^{2}(R_{\text{HMT}})$ values are computed by comparing the reconstructed experimental spectrum with different theoretical curves. In the energy range of interest, the $\chi^{2}(R_{\text{HMT}})$ distribution is shown as red dots with error bars in Fig. 31. A quadratic fit within the range of $5\%<R_{\text{HMT}}<35\%$ to this distribution, represented by the dashed line, reveals that the minimum occurs at $R_{\text{HMT}}=(20.8\pm 0.2)\%$, with the uncertainty representing the fitting error.
Figure 31: $\chi^{2}(R_{\text{HMT}})$ values for different $R_{\text{HMT}}$ values, along with the corresponding quadratic fit, in the energy range from $35~\rm MeV$ to $100~\rm MeV$.
The errors of each bin in the reconstructed $\gamma$ energy spectrum from experimental measurements by RL algorithm approximately follow a Gaussian distribution. Therefore, we can evaluate the confidence intervals for $R_{\text{HMT}}$ based on the $\chi^{2}$ distribution with one degree of freedom (one parameter $R_{\text{HMT}}$). Specifically, the values of $\chi^{2}$ that exceed the minimum by 1, 9 and 25 correspond to confidence levels of $\pm 1\sigma$, $\pm 3\sigma$ and $\pm 5\sigma$, respectively, as indicated by the black dashed lines in the figure. Therefore, the optimal value of $R_{\text{HMT}}$ determined from the reconstructed spectrum is
$R_{\text{HMT}}=(20.8\pm 1.8)\%,$ (36)
This result is consistent with the forward analysis given by Eq. (26) within the error margin.
VII Conclusion and Outlook
In conclusion, SRC in atomic nuclei has drawn increasing attention in nuclear physics, since it is not only relavant to the nuclear structure and the fundamental implication of the nuclear force, but also essential to understand the dense nuclear matter existing in nuclear stars. Moerover, SRC is also related to modification of the parton distribution of nucleons in atomic nuclei (nPDF) according to the QCD factorization scheme. While enormous progress has been made in electron-nucleus scattering experiment and nucleus-proton knock-out reactions, experimental progress in HICs is still scarce.
The bremsstrahlung $\gamma$ emissions in HICs has been identified as a novel probe to detect the SRC because of the fact that the HMT of the nucleons arising from SRC hardens the high energy spectrum of the bremsstrahlung $\gamma$-rays. Following this suggestions, two experiments has been performed on CSHINE. The previous experiment of 86Kr+124Sn at 25 MeV/u suggested the existence of HMT, offering the motivation to detect the SRC using bremsstrahlung $\gamma$ emissions.
To be more quantitative and more precise, the new experiment of 124Sn+124Sn at 25 MeV/u has been performed with much improvements on the experimental setup as well as data analysis, including the installation of the veto scintillators surrounding the CsI(Tl) hodoscope and the extension of the energy range compared to the previous one, as well as the test of the non-linearity of the response of the CsT(Tl) crystal using the quasi monochromatic $\gamma$ beam at SLEGS. The total statistics has been improved. Moreover, the adoption of a symmetric system in the experiment allows to determine the $R_{\rm HMT}$ to the specific nucleus 124Sn.
Two approaches have been applied to obtain the bremsstrahlung $\gamma$-ray energy spectrum. The main approach is to use the whole data set in a slow coincidence gate by subtracting the cosmic-ray muon background which has been measured with beam-off. The other one as a cross check is to use the data in a fast coincidence gate, while the background is obtained in a beam-on random coincidence time window with the same width. With the second approach, the SSD M2  trigger condition is necessarily excluded since the trigger signal suffers seriously time jitter. Both approaches reach the consistent constraint on the HMT fraction within uncertainty. Combined, the fraction of high-momentum nucleons arising from SRC has been derived as $R_{\rm HMT}=(20\pm 3)\%$.
From the experimentally measured spectrum, we also reconstructed the original bremsstrahlung $\gamma$-ray energy spectrum by applying the Richardson-Lucy deblurring algorithm, which is originated from the application of optic deblurring method to solve the inverse problem. The inference of the fraction of the HMT nucleons is in accordance with the above result. Advantageously, the reconstruction of the original $\gamma$ spectrum allows for direct comparison to theoretic calculations without knowing the detector response matrix by other theoretic groups in the future.
The current experimental result gives a precise fraction of SRC nucleons in nuclei, confirming the nonzero $R_{\rm HMT}$ beyond $5\sigma$ confidential level. For the first time, our work establishes a new paradigm to probe the SRC and the parton dynamics in nuclei using low-energy heavy-ion collisions.
Acknowledgement  This work is supported by the National Natural Science Foundation of China under Grant Nos. 12335008, 12205160, 12275129, by the Ministry of Science and Technology under Grant No. 2022YFE0103400, by the Center for High Performance Computing and Initiative Scientific Research Program in Tsinghua University and by HIRFL. The authors acknowledge the HIRFL machine team for providing the excellent beam condition.